[Hallc_running] [New Logentry] Monday RC update

brads at jlab.org brads at jlab.org
Mon Jul 1 14:10:02 EDT 2019

Logentry Text:
Started taking data after the change to 4-pass yesterday evening.   Beam checkout, Heep checks, and some initial 'large-angle' data were taken with reasonable Accelerator uptime over Owl shift.

Beam went away this morning so MCC could do some needed work in the injector, and took advantage of the time to move the SHMS to minimum angle (5.75deg).  Beam is just coming back into the Hall now (2pm).

We have about 5 beam-hours of data to take with the SHMS at this minimum angle in late Day shift and into Swing.  After we've completed that program (mid-Swing), we will take another access to rotate the SHMS out to a nominal 7 degrees where we can remotely control spectrometer rotation again.

There *may* be another short beam studies Tuesday morning, or it could be cancelled.  We'll see how things run this evening.  NB. the 'missing current at the injector' problem seems to be due to a misalignment of the FC2 -- easy fix, and good news.

** NOTE:  The RC meeting today will immediately follow the Hall C meeting at 3pm today.  I will set up Bluejeans using the 'RC' access information before the Hall C meeting starts, so dial in anytime. **


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3697243
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190701/dae8fcf5/attachment.html>

More information about the Hallc_running mailing list